OpenStudy (anonymous):

How do I even solve this?!!! 1/t+1/3t+1/5t=9

OpenStudy (anonymous):

15+5+3=9t 23=9t t=23/9

OpenStudy (espex):

I would give them all the same denominator, multiplying by 15/15 giving you \[15/15t + 5/15t + 3/15t = 9\] Adding them together would give you \[23/15t=9\] Then multiply out your 15t \[23=135t\] and \[t=23/135\]